If x,y €[0,10] then find no of solutions of
 
 3sec x . Sec x -1 . Sqrt (9y2 -6y +2) <= 1

Grade:11

1 Answers

Arun
5 years ago
 
min value of 9y^2 – 6y + 2 =  – (36 – 72) /36 = 1
 
hence
 
now
 
sec x has minimum value = 1
 
hence if we talk about using min value then
 
3^(1 – 1)  . sqrt (1)
 
1 * 1 = 1
 
1
 
hence
secx = 1 and y = 6/18 = 1/3
 
hence
 
x should be equal to n*pi
and y should be equal to 1/3